OpenStudy (anonymous):

use DeMoivre's theorem to simplify the expression. write the answer in a + b form. (1- i)^10. pleeeaaaasssseee help!!!

OpenStudy (anonymous):

(e^j(-pi/4))^10

OpenStudy (anonymous):

i don't understand please

OpenStudy (anonymous):

r = sqrt(2) theta = -pi/4

OpenStudy (anonymous):

\[(1-j)^{10} = \sqrt{2}^{10} * (\cos10(-\pi/4) + jsin(10(-\pi/4))\]

OpenStudy (anonymous):

woow u are too awsome..but i need to write the answer in a + bi form

OpenStudy (anonymous):

you can convert from polar to rectangular

OpenStudy (anonymous):

how do i do that please?

OpenStudy (anonymous):

Thanks for the help I am going to get my mom to check my work I want to understand what I am doing so I can learn math not just answer questions for the teacher. Thank you all again.